OpenStudy (anonymous):

Find the linear approximation of the tangent line to f(x)=1/the square root of 1-x at x=0

OpenStudy (anonymous):

-1/2 i think

OpenStudy (anonymous):

f(x) = 1/sqrt(1-x) f'(x) = -1/2sqrt(1-x) f'(0) = -1/2

OpenStudy (anonymous):

Linear approximation around x = x0 in general: \[y - f(x_{0}) = f'(x_{0})(x - x_{0})\] \[y= f'(x_{0})(x - x_{0}) +f(x_{0})\] \[f(x) = 1/\sqrt{1-x}\] \[f'(x) = 1/(2\sqrt{(1 - x)^3})\] so our approximation: \[f(0) = 1\] \[f'(0) = 1/2\] \[y = (1/2)(x - 0) + 1\] \[y = (1/2)x + 1\]

OpenStudy (anonymous):

oh man...i got that basic derivative wrong...sorry...ignore my nonsensical post